As sales manager, Joe Batista was given the following static budget report for selling expenses in the Clothing Department of Soria Company for the month of October.

SORIA COMPANY
Clothing Department
Budget Report
For the Month Ended October 31, 2020





Difference


Budget


Actual

Favorable
Unfavorable
Neither Favorable
nor Unfavorable
Sales in units
7,900

11,000

3,100
Favorable
Variable expenses






    Sales commissions
$2,054

$2,860

$806
Unfavorable
    Advertising expense
869

770

99
Favorable
    Travel expense
3,476

4,950

1,474
Unfavorable
    Free samples given out
1,659

1,210

449
Favorable
       Total variable
8,058

9,790

1,732
Unfavorable
Fixed expenses






     Rent
1,900

1,900

–0–
Neither Favorable nor Unfavorable
     Sales salaries
1,100

1,100

–0–
Neither Favorable nor Unfavorable
     Office salaries
800

800

–0–
Neither Favorable nor Unfavorable
     Depreciation—autos (sales staff)
600

600

–0–
Neither Favorable nor Unfavorable
       Total fixed
4,400

4,400

–0–
Neither Favorable nor Unfavorable
Total expenses
$12,458

$14,190

$1,732
Unfavorable

As a result of this budget report, Joe was called into the president’s office and congratulated on his fine sales performance. He was reprimanded, however, for allowing his costs to get out of control. Joe knew something was wrong with the performance report that he had been given. However, he was not sure what to do, and comes to you for advice.

Prepare a budget report based on flexible budget data to help Joe. (List variable costs before fixed costs.)
Answer #1
Budget Actual Difference
Sales in units 11000 11000
Variable expenses
Sales commissions 2860 2860 0 Neither Favorable nor Unfavorable
Advertising expense 1210 770 440 Favorable
Travel expense 4840 4950 110 Unfavorable
Free samples 2310 1210 1100 Favorable
Total variable 11220 9790 1430 Favorable
Fixed expenses
Rent 1900 1900 0 Neither Favorable nor Unfavorable
Sales salaries 1100 1100 0 Neither Favorable nor Unfavorable
Office salaries 800 800 0 Neither Favorable nor Unfavorable
Depreciation—autos (sales staff) 600 600 0 Neither Favorable nor Unfavorable
Total fixed 4400 4400 0 Neither Favorable nor Unfavorable
Total expenses 15620 14190 1430 Favorable
Workings:
Variable expenses
Sales commissions 2860 =2054/7900*11000
Advertising expense 1210 =869/7900*11000
Travel expense 4840 =3476/7900*11000
Free samples 2310 =1659/7900*11000